sam is 4 years older than his wife and 31 years older than his son the ages add up to 82 years. Let Sam be x years. Write an equation and solve it to find x and the age of Sam's wife and son

Answer by rothauserc(4718)

Sam's age is x, Sam's wife's age is x-4, and Sam's son's age is x-31, then
:
x +x-4 +x-31 = 82
:
3x -35 = 82
:
3x = 117
:
x = 39
:
****************************************
Sam's wife is 39 -4 = 35 years old
:
Sam's son's age is 39 -31 = 8 years old
:
check answer
:
39 +35 +8 = 82
:
82 = 82
:
answer checks
*****************************************
: